# decoding inner speech with an end-to-end brain-to-text neural interface

### v2

@@ 5.2 Ablating large language model decoder
We conduct an ablation study to quantify the effects of LLM decoder type and model design on attempted speech decoding performance. Specifically, we compare text-based and audio-based LLMs, and test different prompt designs by treating neural embeddings as either audio or neural modalities (Section 3.2 and Figure 3B). Finally, we evaluate how contrastive learning for cross-modal alignment influences decoding performance. We focus on models with 1 $\sim$ 7B parameters, which strike a balance between expressive capacity and adaptability for neural decoding with scarce labeled data. For text-based LLMs, we use Qwen2.5-1.5B, Qwen2.5-7B (qwen2025qwen25technicalreport), Qwen3-0.6B, and Qwen3-1.7B (yang2025qwen3). For audio-LLMs, we use Aero1-Audio 1.5B (li2025aero), an audio-extended version of Qwen2.5-1.5B, and Qwen2-Audio-7B (chu2024qwen2), a larger model capturing richer acoustic representations. Comparing these audio-extended models with their text-only counterparts allows us to isolate the impact of audio processing from general language modeling capacity.
@@
Our experiments (Figure 3C–D) show that, for models of comparable size, audio-based LLMs outperform text-based LLMs in decoding, with Aero1-Audio 1.5B (li2025aero), the audio-extended variant of Qwen2.5 1.5B, consistently achieving the best results. This suggests that pretraining LLMs on audio tasks provides inductive biases closer to the neural decoding problem, enabling a shallow MLP to align modality more effectively. Treating neural encoder outputs as a neural modality performs slightly better than treating them as audio, indicating that while neural embeddings need not be interpreted as audio, LLMs still benefit from speech knowledge. We also find that smaller LLMs generally achieve lower WERs than larger LLMs when trained with limited labeled examples, since speech BCI tasks require English translation rather than advanced reasoning, which is typically seen in models above 7B parameters. Finally, using contrastive learning to align neural and text embeddings further improves performance.
@@ 5.3 Aligning neural embeddings for cross-task generalization
Although BIT is trained for speech decoding, we want to understand what semantic content is encoded in its latent embeddings. To this end, we conduct sentence-level representational similarity analysis (RSA) (kriegeskorte2008representational) to compare the representational geometry of neural embeddings against that of LLMs. If neural embeddings preserve relational structures similar to LLMs, they may provide more informative features for decoding. Importantly, all neural encoders in the RSA analysis are taken prior to sentence-level fine-tuning to ensure that we compare the intrinsic representational geometry rather than the effect of supervised adaptation (details in Appendix L.1). Figure 4A shows the RSA scores between neural and audio-LLM text embeddings. Pretrained encoder outputs are more similar to audio-LLM representations than those of RNN and BIT-TFS , indicating that large-scale pretraining facilitates linguistically structured representations.
@@ 6 Discussion
Limitations and future directions: Our approach faces several challenges and opportunities for improvement. First, the end-to-end approach requires about 0.95 seconds per sentence on average, which is slower than the cascaded approach (0.24 seconds) and inefficient for real-time BCIs. In addition, the neural encoder is trained with bidirectional attention to maximize performance, rendering it unsuitable for online decoding; switching to causal attention is feasible, albeit at some cost to decoding accuracy. Similarly, while we employ a compact 1.5B audio-LLM, larger models cannot run on-device, further limiting real-time applications. Another consideration is pretraining: we find that human data yield larger gains than cross-species transfer, likely because human datasets include tasks more relevant to speech (and, to a lesser extent, handwriting), whereas monkey reaching data are less applicable. Additionally, the neural encoder requires substantial unlabeled datasets to address sensor variability, and LLMs need large amounts of labeled data to outperform cascaded approaches. However, we have limited access to private human data. Finally, the LLM decoder itself can be improved through better modality alignment and prompt design, and there remains a broader need to develop methods addressing non-stationarity (le2025spint), neural plasticity, and user-interface co-adaptation for long-term BCI use.
Broader impact: This work improves current cascaded speech BCI systems to restore high-accuracy communication for users with paralysis. It also advances SOTA end-to-end models, making them easier to optimize and deploy in real-time applications. These developments could enable patients who are unable to communicate or move to interact with the outside world and perform everyday tasks with AI assistance. At the same time, it raises important safety and privacy concerns, as decoding inner speech must require explicit user consent and reliable safeguards (kunz2025inner).
@@ A.1 Human data
karpowicz2024few.
This dataset is from the Few-shot Algorithms for Consistent Neural Decoding (FALCON) Benchmark, which evaluates intracortical BCI decoders for stability across sessions. We use only the human handwriting subset, also associated with fan2023plug, available at DANDI. The threshold spike data in this subset also include 192 electrodes.
card2024accurate.
@@ Appendix B Metrics
We use a 41-token phoneme vocabulary, including all phonemes plus silence and the CTC blank token. The phoneme error count is first computed by removing repeats and blanks from the decoded phoneme sequences, then calculating the Levenshtein edit distance (levenshtein1966binary) to the ground-truth sequences, counting substitutions, insertions, and deletions. The phoneme error rate is then obtained by normalizing the error count by the total number of ground-truth phonemes.
@@ Appendix C Phoneme-level decoding
In the end-to-end model, phonemes serve as intermediate targets that encode phonetic information into neural representations to guide LLM fine-tuning, even when the final goal is sentence prediction. We observe that, without phoneme-decoding training, simply pretraining the neural encoder on a large data corpus does not lead to low WERs during the sentence-level decoding stage, highlighting the importance of phonetic information for speech decoding tasks. This phenomenon is also reported in previous end-to-end methods (feng2024towards), which employ a two-stage decoding strategy that trains the model for phoneme prediction before sentence prediction.
@@ Appendix D Cascaded decoder
The cascaded decoder consists of a 5-gram LM and an OPT model for re-scoring the LM outputs. In this section, we briefly describe the resources and procedures used; for comprehensive details, readers are referred to the original studies (willett2023high; card2024accurate).
@@
We use the 5-gram LM provided by the Brain-to-Text Benchmark ’24 and ’25 (willett2023high; card2024accurate) to decode word sequences from neural encoder outputs. The pretrained model is publicly available on DRYAD. The 5-gram LM is created using OpenWebText2 corpus and therefore has a large vocabulary, constructed from 634 million sentences with 99 billion words. The LM decoder performs an approximate Viterbi (beam) search, which maps CTC phoneme label sequences from the encoder to words. During inference, the beam search integrates state transition probabilities with the encoder’s predicted CTC phoneme label probabilities to find the most likely word sequences. See Section 8 of the supplementary materials for details on constructing the 5-gram LM and performing inference with it. In Table 5, we report the parameters used for beam search with a 5-gram LM.
@@
Given the sentence hypotheses from the 5-gram LM, an OPT model (zhang2022opt) is employed to rescore the candidate sentences and further improve decoding accuracy. A publicly available pretrained OPT with 6.7B parameters is used to rescore the n-best outputs from the n-gram LM. The re-scoring procedure considers both the probability of the sentence’s corresponding CTC phoneme label sequence output by the encoder (acoustic score) and the sentence probability estimated by the n-gram LM; see Section 8.5 of the supplementary materials for details on the equation used for OPT re-scoring and the LM decoding parameters. In Table 5, we report the parameters used for language model re-scoring with OPT.
@@ Appendix E End-to-End Decoder
We include models from the Qwen2.5 and Qwen3 families. Qwen2.5-1.5B (qwen2025qwen25technicalreport) is a mid-sized LLM trained with an improved pretraining pipeline and instruction tuning. We also include Qwen2.5-7B. To examine the effect of model size and architectural advances, we include two recent models from the Qwen3 series: Qwen3-0.6B (yang2025qwen3), a compact model optimized for efficiency, and Qwen3-1.7B, a larger variant designed to provide stronger language modeling capacity while remaining computationally tractable. Together, these models provide a representative set of text-based LLMs at approximately 1B scale, enabling comparison with audio-augmented counterparts under similar parameter regimes.
@@
In addition to text-only baselines, we consider audio-augmented LLMs. Aero1-Audio 1.5B (li2025aero) extends Qwen2.5-1.5B with an audio front-end, enabling the model to process acoustic representations alongside text. To understand the effect of model scale in the audio domain, we also include Qwen2-Audio 7B (chu2024qwen2), a larger audio-based LLM capturing richer acoustic and semantic features. Comparing Aero1-Audio 1.5B and Qwen2-Audio 7B with text-only models of similar size allows us to quantify the contributions of audio conditioning relative to purely textual modeling in speech decoding from neural activity.
@@ Appendix G Model ensemlbing using LLM merger
To further improve the decoding accuracy of the cascaded and end-to-end decoders for benchmark submissions, we employ a model ensembling approach with LLM merging (li2024brain). While the 5-gram and OPT models can partially correct phoneme decoding errors, their outputs are not always perfect, and different neural encoders produce varying sentence candidates. To mitigate these errors, a fine-tuned LLM can be used to evaluate multiple sentence candidates from an ensemble of cascaded or end-to-end encoders. Using both the transcription candidates and their corresponding phoneme sequences as inputs, li2024brain finetuned a GPT-3.5 to generate the correct sentence. This setup enables the model to learn the relationship between predicted phonemes and decoded sentences, as well as to identify common, model-specific errors across decoder outputs. For both cascaded and end-to-end approaches, we provide the candidate sentences along with the phoneme outputs from neural encoders that are randomly initialized with different seeds to the fine-tuned GPT-3.5, which then generates the final best sentence used for evaluation.
@@ Appendix H Brain-to-Text Benchmark ’24 and ’25 Details
The top scores on the leaderboards for both cascaded and end-to-end methods are achieved using model ensembling with LLM merging. For this benchmark, we initialize ten cross-species, cross-task pretrained encoders with different random seeds, each fine-tuned to decode sentences from participant T12 data. Each model is trained on the training data provided by the benchmark and uses the provided validation set for checkpoint selection. Each of the ten models outputs its own phoneme and sentence sequences, which are then provided to the fine-tuned GPT-3.5 to select the final best sentence sequence. This procedure is applied to both the entries BIT Cascaded + Ensemble and BIT End-to-End + Ensemble , which are the top-ranked submissions for the cascaded and end-to-end approaches on the leaderboard. All of our other leaderboard entries are based on a single model with a single seed. For information on the other leaderboard entries, see willett2024brain.
@@
The top scores on the leaderboards for both cascaded and end-to-end methods are achieved using model ensembling with LLM merging. For this benchmark, we initialize ten cross-species, cross-task pretrained encoders with different random seeds, each fine-tuned to decode sentences from participant T15 data. The LLM merging is therefore based on the phoneme and sentence outputs from these ten models. All of our other leaderboard entries are based on a single model with a single seed. As the benchmark is ongoing at the time of this submission, we only report the provided baselines on the leaderboard. In Table 2, the RNN (baseline) uses a single RNN encoder with a 5-gram LM, labeled “UCD-NPL causal RNN + 5gram.” The RNN + Ensemble combines multiple RNNs with LLM merging, labeled “Stanford-NPTL causal RNN Ensemble + 5gram.” The RNN-TTA + Pseudo-Ensemble generates diverse outputs from a single RNN model via noise injection and selects the final sentence with an LLM merger; it is listed as “Stanford-NPTL causal RNN TTA-Ensemble + 5gram,” where TTA stands for test-time adaptation.
@@ Appendix I Contrastive learning for modality alignment
Contrastive learning is then applied to pull together corresponding neural and text modality tokens (positive pairs) while pushing apart tokens from different samples (negative pairs) within a training batch. Following a symmetric InfoNCE loss (oord2018representation; radford2021learning), the loss for a batch is defined as
@@
We restrict the analysis to sentences with neural time-patch token lengths between 45 $\sim$ 80 for participant T12 and 120 $\sim$ 200 for participant T15, ensuring sufficient temporal resolution and balanced coverage (see Figure 5). Neural embeddings are converted into fixed-length vectors using segmented mean pooling, where each sequence is divided into 10 segments and the segment averages are concatenated to preserve coarse temporal dynamics. In contrast, LLM embeddings are reduced to sentence vectors via mean pooling across tokens. For each modality (neural and text), we compute a representational dissimilarity matrix (RDM) (kriegeskorte2008representational) using one minus the cosine similarity. To compare representational structures, we extract the upper-triangular entries of each RDM and compute the Pearson correlation coefficient between neural and LLM RDMs. The resulting RSA score quantifies how well the geometry of neural embeddings aligns with language structures in LLMs. Framing RSA as an interpretability tool, this analysis demonstrates that the transformer encoder learns representations more consistent with language-level structure than RNNs. RSA is applied only to attempted speech data from participants T12 and T15 in the studies (willett2023high; card2024accurate).
@@
We fit an LDA using scikit-learn. For each word, neural embeddings are first projected onto the PCA subspace and then averaged across time and trials to produce a single feature vector. The resulting low-dimensional neural projections are used as features to fit the LDA to predict trial type (attempted vs. imagined speech for binary classification). We restrict the LDA to a single dimension to identify an axis that linearly separates neural projections from the two tasks for visualization purposes. However, when computing Euclidean distances between attempted and imagined speech embeddings, we use the original high-dimensional embeddings for each word before PCA projection to ensure comparisons reflect both direction and magnitude in the embedding space. These analyses are applied only to attempted and imagined speech data from participants T12 and T15 in the study (kunz2025inner).
@@
We implement a transformer encoder with an architecture similar to feghhi2025time. Since their official code is not publicly available at the time of this submission, we construct our own architecture using custom model parameters listed in Table 10. To select the optimal model architecture and optimizer configuration for training on single-subject data (T12 and T15), we use Ray Tune (liaw2018tune) to randomly sample 30 optimizer hyperparameter (batch size, weight decay, and learning rate) combinations from the ranges listed in Table 12. The optimizer hyperparameters for single-subject phoneme-level decoding is selected based on these ranges. For imagined speech decoding, we reuse the same model hyperparameters for both participants. For pretraining, we reuse the same model architecture while adjusting the learning rate, weight decay and batch size to $5\times 10^{-4}$, $1\times 10^{-5}$ and 64, respectively. We do not increase the model depth when scaling to human and monkey datasets, as SSL pretraining performance (evaluated using the validation $R^{2}$ metric for neural activity reconstruction) shows no decline, indicating that the model capacity is sufficient to handle the data variability. The transformer model comprises $\sim$ 7 million parameters. The inclusion of subject-specific patch embedding modules and linear decoders increases the total number of model parameters to $\sim$ 13 million.
@@
For the RNN encoder, we use the PyTorch implementation provided by the benchmark host in the official GitHub repository. We retain the original model architecture (Table 11) and apply Ray Tune (liaw2018tune) to randomly sample 30 optimizer hyperparameter (batch size, weight decay, and learning rate) combinations from the ranges listed in Table 12, using attempted speech data from participants T12 and T15. For imagined speech decoding, we reuse the same model hyperparameters for both participants, except for the number of day-specific read-in layers, which differs across datasets.
@@
For both pretraining and fine-tuning, we train the model using the AdamW optimizer (loshchilov2017decoupled). Encoders pretrained on human-only or human-and-monkey datasets are trained on a single NVIDIA A100 GPU (80 GB memory) in under 2 days for a total of 400 epochs. Phoneme decoding is performed on a single NVIDIA A40 (48 GB memory) or A100 GPU (40 or 80 GB memory) in less than 8 hours for T12 and less than 1 day for T15, for 800 epochs. For hyperparameter tuning, we use 4 NVIDIA A40 GPUs for phoneme decoding, completing the process in less than 2 days. During training, we use validation metrics ($R^{2}$, PER) to select the best model checkpoint, based on the highest $R^{2}$ or lowest PER. We employ mixed precision (bfloat16) to improve training efficiency for both pretraining and fine-tuning, and all models fit on a single NVIDIA GPU.

### v3

@@ 5.2 Ablating large language model decoder
We conduct an ablation study to quantify the effects of LLM decoder type and model design on attempted speech decoding performance. Specifically, we compare text-based and audio-based LLMs, and test different prompt designs by treating neural embeddings as either audio or neural modalities (Section 3.2 and Figure 3B). Finally, we evaluate how contrastive learning for cross-modal alignment influences decoding performance. We focus on models with 1 $\sim$ 7B parameters, which strike a balance between expressive capacity and adaptability for neural decoding with scarce labeled data. For text-based LLMs, we use Qwen2.5-1.5B, Qwen2.5-7B (Qwen et al., 2025), Qwen3-0.6B, and Qwen3-1.7B (Yang et al., 2025). For audio-LLMs, we use Aero1-Audio 1.5B (Li et al., 2025), an audio-extended version of Qwen2.5-1.5B, and Qwen2-Audio-7B (Chu et al., 2024), a larger model capturing richer acoustic representations. Comparing these audio-extended models with their text-only counterparts allows us to isolate the impact of audio processing from general language modeling capacity.
@@
Our experiments (Figure 3C–D) show that, for models of comparable size, audio-based LLMs outperform text-based LLMs in decoding, with Aero1-Audio 1.5B (Li et al., 2025), the audio-extended variant of Qwen2.5 1.5B, consistently achieving the best results. This suggests that pretraining LLMs on audio tasks provides inductive biases closer to the neural decoding problem, enabling a shallow MLP to align modality more effectively. Treating neural encoder outputs as a neural modality performs slightly better than treating them as audio, indicating that while neural embeddings need not be interpreted as audio, LLMs still benefit from speech knowledge. We also find that smaller LLMs generally achieve lower WERs than larger LLMs when trained with limited labeled examples, since speech BCI tasks require English translation rather than advanced reasoning, which is typically seen in models above 7B parameters. Finally, using contrastive learning to align neural and text embeddings further improves performance.
@@ 5.3 Aligning neural embeddings for cross-task generalization
Although BIT is trained for speech decoding, we want to understand what semantic content is encoded in its latent embeddings. To this end, we conduct sentence-level representational similarity analysis (RSA) (Kriegeskorte et al., 2008) to compare the representational geometry of neural embeddings against that of LLMs. If neural embeddings preserve relational structures similar to LLMs, they may provide more informative features for decoding. Importantly, all neural encoders in the RSA analysis are taken prior to sentence-level fine-tuning to ensure that we compare the intrinsic representational geometry rather than the effect of supervised adaptation (details in Appendix L.1). Figure 4A shows the RSA scores between neural and audio-LLM text embeddings. Pretrained encoder outputs are more similar to audio-LLM representations than those of RNN and BIT-TFS , indicating that large-scale pretraining facilitates linguistically structured representations.
@@ 6 Discussion
Limitations and future directions: Our approach faces several challenges and opportunities for improvement. First, the end-to-end approach requires about 0.95 seconds per sentence on average, which is slower than the cascaded approach (0.24 seconds) and inefficient for real-time BCIs. In addition, the neural encoder is trained with bidirectional attention to maximize performance, rendering it unsuitable for online decoding; switching to causal attention is feasible, albeit at some cost to decoding accuracy. Similarly, while we employ a compact 1.5B audio-LLM, larger models cannot run on-device, further limiting real-time applications. Another consideration is pretraining: we find that human data yield larger gains than cross-species transfer, likely because human datasets include tasks more relevant to speech (and, to a lesser extent, handwriting), whereas monkey reaching data are less applicable. Additionally, the neural encoder requires substantial unlabeled datasets to address sensor variability, and LLMs need large amounts of labeled data to outperform cascaded approaches. However, we have limited access to private human data. Finally, the LLM decoder itself can be improved through better modality alignment and prompt design, and there remains a broader need to develop methods addressing non-stationarity (Le et al., 2025), neural plasticity, and user-interface co-adaptation for long-term BCI use.
Broader impact: This work improves current cascaded speech BCI systems to restore high-accuracy communication for users with paralysis. It also advances SOTA end-to-end models, making them easier to optimize and deploy in real-time applications. These developments could enable patients who are unable to communicate or move to interact with the outside world and perform everyday tasks with AI assistance. At the same time, it raises important safety and privacy concerns, as decoding inner speech must require explicit user consent and reliable safeguards (Kunz et al., 2025).
@@ A.1 Human data
Karpowicz et al. (2024).
This dataset is from the Few-shot Algorithms for Consistent Neural Decoding (FALCON) Benchmark, which evaluates intracortical BCI decoders for stability across sessions. We use only the human handwriting subset, also associated with Fan et al. (2023), available at DANDI. The threshold spike data in this subset also include 192 electrodes.
Card et al. (2024).
@@ Appendix B Metrics
We use a 41-token phoneme vocabulary, including all phonemes plus silence and the CTC blank token. The phoneme error count is first computed by removing repeats and blanks from the decoded phoneme sequences, then calculating the Levenshtein edit distance (Levenshtein, 1966) to the ground-truth sequences, counting substitutions, insertions, and deletions. The phoneme error rate is then obtained by normalizing the error count by the total number of ground-truth phonemes.
@@ Appendix C Phoneme-level decoding
In the end-to-end model, phonemes serve as intermediate targets that encode phonetic information into neural representations to guide LLM fine-tuning, even when the final goal is sentence prediction. We observe that, without phoneme-decoding training, simply pretraining the neural encoder on a large data corpus does not lead to low WERs during the sentence-level decoding stage, highlighting the importance of phonetic information for speech decoding tasks. This phenomenon is also reported in previous end-to-end methods (Feng et al., 2024), which employ a two-stage decoding strategy that trains the model for phoneme prediction before sentence prediction.
@@ Appendix D Cascaded decoder
The cascaded decoder consists of a 5-gram LM and an OPT model for re-scoring the LM outputs. In this section, we briefly describe the resources and procedures used; for comprehensive details, readers are referred to the original studies (Willett et al., 2023; Card et al., 2024).
@@
We use the 5-gram LM provided by the Brain-to-Text Benchmark ’24 and ’25 (Willett et al., 2023; Card et al., 2024) to decode word sequences from neural encoder outputs. The pretrained model is publicly available on DRYAD. The 5-gram LM is created using OpenWebText2 corpus and therefore has a large vocabulary, constructed from 634 million sentences with 99 billion words. The LM decoder performs an approximate Viterbi (beam) search, which maps CTC phoneme label sequences from the encoder to words. During inference, the beam search integrates state transition probabilities with the encoder’s predicted CTC phoneme label probabilities to find the most likely word sequences. See Section 8 of the supplementary materials for details on constructing the 5-gram LM and performing inference with it. In Table 5, we report the parameters used for beam search with a 5-gram LM.
@@
Given the sentence hypotheses from the 5-gram LM, an OPT model (Zhang et al., 2022) is employed to rescore the candidate sentences and further improve decoding accuracy. A publicly available pretrained OPT with 6.7B parameters is used to rescore the n-best outputs from the n-gram LM. The re-scoring procedure considers both the probability of the sentence’s corresponding CTC phoneme label sequence output by the encoder (acoustic score) and the sentence probability estimated by the n-gram LM; see Section 8.5 of the supplementary materials for details on the equation used for OPT re-scoring and the LM decoding parameters. In Table 5, we report the parameters used for language model re-scoring with OPT.
@@ Appendix E End-to-End Decoder
We include models from the Qwen2.5 and Qwen3 families. Qwen2.5-1.5B (Qwen et al., 2025) is a mid-sized LLM trained with an improved pretraining pipeline and instruction tuning. We also include Qwen2.5-7B. To examine the effect of model size and architectural advances, we include two recent models from the Qwen3 series: Qwen3-0.6B (Yang et al., 2025), a compact model optimized for efficiency, and Qwen3-1.7B, a larger variant designed to provide stronger language modeling capacity while remaining computationally tractable. Together, these models provide a representative set of text-based LLMs at approximately 1B scale, enabling comparison with audio-augmented counterparts under similar parameter regimes.
@@
In addition to text-only baselines, we consider audio-augmented LLMs. Aero1-Audio 1.5B (Li et al., 2025) extends Qwen2.5-1.5B with an audio front-end, enabling the model to process acoustic representations alongside text. To understand the effect of model scale in the audio domain, we also include Qwen2-Audio 7B (Chu et al., 2024), a larger audio-based LLM capturing richer acoustic and semantic features. Comparing Aero1-Audio 1.5B and Qwen2-Audio 7B with text-only models of similar size allows us to quantify the contributions of audio conditioning relative to purely textual modeling in speech decoding from neural activity.
@@ Appendix G Model ensemlbing using LLM merger
To further improve the decoding accuracy of the cascaded and end-to-end decoders for benchmark submissions, we employ a model ensembling approach with LLM merging (Li et al., 2024). While the 5-gram and OPT models can partially correct phoneme decoding errors, their outputs are not always perfect, and different neural encoders produce varying sentence candidates. To mitigate these errors, a fine-tuned LLM can be used to evaluate multiple sentence candidates from an ensemble of cascaded or end-to-end encoders. Using both the transcription candidates and their corresponding phoneme sequences as inputs, Li et al. (2024) finetuned an LLM to generate the correct sentence. This setup enables the model to learn the relationship between predicted phonemes and decoded sentences, as well as to identify common, model-specific errors across decoder outputs. For both cascaded and end-to-end approaches, we provide the candidate sentences along with the phoneme outputs from neural encoders that are randomly initialized with different seeds to the fine-tuned LLM, which then generates the final best sentence used for evaluation.
@@ Appendix H Brain-to-Text Benchmark ’24 and ’25 Details
The top scores on the leaderboards for both cascaded and end-to-end methods are achieved using model ensembling with LLM merging. For this benchmark, we initialize 10 cross-species, cross-task pretrained encoders with different random seeds, each fine-tuned to decode sentences from participant T12 data. Each model is trained on the training data provided by the benchmark and uses the provided validation set for checkpoint selection. Each of the 10 models outputs its own phoneme and sentence sequences, which are then provided to a fine-tuned GPT-3.5 to select the final best sentence sequence. This procedure is applied to both the entries BIT Cascaded + Ensemble and BIT End-to-End + Ensemble , which are the top-ranked submissions for the cascaded and end-to-end approaches on the leaderboard. All of our other leaderboard entries are based on a single model with a single seed. For information on the other leaderboard entries, see Willett et al. (2024).
@@
The top scores on the public leaderboards for both cascaded and end-to-end methods are achieved using model ensembling with LLM merging. For this benchmark, we initialize 29 cross-species, cross-task pretrained encoders with different random seeds, each fine-tuned to decode sentences from participant T15 data. We fine-tune GPT-4 to merge the phoneme and sentence outputs from these 29 models. All of our other leaderboard entries are based on a single model with a single seed. In Table 2, the RNN (baseline) uses a single RNN encoder with a 5-gram LM, labeled “UCD-NPL causal RNN + 5gram.” The RNN + Ensemble combines multiple RNNs with LLM merging, labeled “Stanford-NPTL causal RNN Ensemble + 5gram.” The RNN-TTA + Pseudo-Ensemble generates diverse outputs from a single RNN model via noise injection and selects the final sentence with an LLM merger; it is listed as “Stanford-NPTL causal RNN TTA-Ensemble + 5gram,” where TTA stands for test-time adaptation.
@@ Appendix I Contrastive learning for modality alignment
Contrastive learning is then applied to pull together corresponding neural and text modality tokens (positive pairs) while pushing apart tokens from different samples (negative pairs) within a training batch. Following a symmetric InfoNCE loss (Oord et al., 2018; Radford et al., 2021), the loss for a batch is defined as
@@
We restrict the analysis to sentences with neural time-patch token lengths between 45 $\sim$ 80 for participant T12 and 120 $\sim$ 200 for participant T15, ensuring sufficient temporal resolution and balanced coverage (see Figure 5). Neural embeddings are converted into fixed-length vectors using segmented mean pooling, where each sequence is divided into 10 segments and the segment averages are concatenated to preserve coarse temporal dynamics. In contrast, LLM embeddings are reduced to sentence vectors via mean pooling across tokens. For each modality (neural and text), we compute a representational dissimilarity matrix (RDM) (Kriegeskorte et al., 2008) using one minus the cosine similarity. To compare representational structures, we extract the upper-triangular entries of each RDM and compute the Pearson correlation coefficient between neural and LLM RDMs. The resulting RSA score quantifies how well the geometry of neural embeddings aligns with language structures in LLMs. Framing RSA as an interpretability tool, this analysis demonstrates that the transformer encoder learns representations more consistent with language-level structure than RNNs. RSA is applied only to attempted speech data from participants T12 and T15 in the studies (Willett et al., 2023; Card et al., 2024).
@@
We fit an LDA using scikit-learn. For each word, neural embeddings are first projected onto the PCA subspace and then averaged across time and trials to produce a single feature vector. The resulting low-dimensional neural projections are used as features to fit the LDA to predict trial type (attempted vs. imagined speech for binary classification). We restrict the LDA to a single dimension to identify an axis that linearly separates neural projections from the two tasks for visualization purposes. However, when computing Euclidean distances between attempted and imagined speech embeddings, we use the original high-dimensional embeddings for each word before PCA projection to ensure comparisons reflect both direction and magnitude in the embedding space. These analyses are applied only to attempted and imagined speech data from participants T12 and T15 in the study (Kunz et al., 2025).
@@
We implement a transformer encoder with an architecture similar to Feghhi et al. (2025). Since their official code is not publicly available at the time of this submission, we construct our own architecture using custom model parameters listed in Table 10. To select the optimal model architecture and optimizer configuration for training on single-subject data (T12 and T15), we use Ray Tune (Liaw et al., 2018) to randomly sample 30 optimizer hyperparameter (batch size, weight decay, and learning rate) combinations from the ranges listed in Table 12. The optimizer hyperparameters for single-subject phoneme-level decoding is selected based on these ranges. For imagined speech decoding, we reuse the same model hyperparameters for both participants. For pretraining, we reuse the same model architecture while adjusting the learning rate, weight decay and batch size to $5\times 10^{-4}$, $1\times 10^{-5}$ and 64, respectively. We do not increase the model depth when scaling to human and monkey datasets, as SSL pretraining performance (evaluated using the validation $R^{2}$ metric for neural activity reconstruction) shows no decline, indicating that the model capacity is sufficient to handle the data variability. The transformer model comprises $\sim$ 7 million parameters. The inclusion of subject-specific patch embedding modules and linear decoders increases the total number of model parameters to $\sim$ 13 million.
@@
For the RNN encoder, we use the PyTorch implementation provided by the benchmark host in the official GitHub repository. We retain the original model architecture (Table 11) and apply Ray Tune (Liaw et al., 2018) to randomly sample 30 optimizer hyperparameter (batch size, weight decay, and learning rate) combinations from the ranges listed in Table 12, using attempted speech data from participants T12 and T15. For imagined speech decoding, we reuse the same model hyperparameters for both participants, except for the number of day-specific read-in layers, which differs across datasets.
@@
For both pretraining and fine-tuning, we train the model using the AdamW optimizer (Loshchilov and Hutter, 2017). Encoders pretrained on human-only or human-and-monkey datasets are trained on a single NVIDIA A100 GPU (80 GB memory) in under 2 days for a total of 400 epochs. Phoneme decoding is performed on a single NVIDIA A40 (48 GB memory) or A100 GPU (40 or 80 GB memory) in less than 8 hours for T12 and less than 1 day for T15, for 800 epochs. For hyperparameter tuning, we use 4 NVIDIA A40 GPUs for phoneme decoding, completing the process in less than 2 days. During training, we use validation metrics ($R^{2}$, PER) to select the best model checkpoint, based on the highest $R^{2}$ or lowest PER. We employ mixed precision (bfloat16) to improve training efficiency for both pretraining and fine-tuning, and all models fit on a single NVIDIA GPU.